OpenStudy (anonymous):

Help, I'm having issues with negatives in this problem, cant get right answer: Y+4/y+5-2 Please help me.

OpenStudy (anonymous):

is it an equation?

OpenStudy (anonymous):

yes, a rational expression type thing

OpenStudy (anonymous):

y+4/y+3 but what is your exercise asking?

OpenStudy (anonymous):

"Perform indicated operation"

OpenStudy (anonymous):

does it give the value of y?

OpenStudy (phi):

is the problem \[ \frac{y+4}{y+5} -2 \] ?

OpenStudy (phi):

If so, put the 2 over a common denominator by multiplying it by \[ \frac{2}{1}\frac{(y+5)}{(y+5)} \]

OpenStudy (phi):

now the problem becomes \[ \frac{y+4 -2(y+5)}{y+5} \]

OpenStudy (phi):

can you finish?

OpenStudy (anonymous):

How do i do the negatives?

OpenStudy (anonymous):

I keep getting the wrong answer :(

OpenStudy (anonymous):

multiplication - + = - -- = +

OpenStudy (anonymous):

y + 4 - 2y - 10 / y + 5

OpenStudy (anonymous):

Can you do that now?

OpenStudy (phi):

mel showed the answer, but let's start with -2(y+5) this means multiply -2 times each term in the parentheses. -2*y + -2*5 or -2y-10 so the top which starts y+4 - 2(y+5) becomes y+4 -2y -10 can you follow?

OpenStudy (anonymous):

yes

OpenStudy (phi):

now combine like terms. we can rearrange it to be y -2y +4 -10 y-2y is the same as 1y-2y . Can you do 1y-2y?

OpenStudy (phi):

can you do 1-2? 1y-2y is the same as (1-2)y

OpenStudy (anonymous):

is it y

OpenStudy (anonymous):

Actually, it is -y. Because 2 > 1.

OpenStudy (phi):

Here is a video on negative numbers http://www.khanacademy.org/math/arithmetic/negative-numbers/v/adding-subtracting-negative-numbers It might help?

OpenStudy (anonymous):

Ok, thank you :)

OpenStudy (phi):

This one might be useful (or not!) http://www.khanacademy.org/math/algebra/ck12-algebra-1/v/variable-expressions But first do the negative numbers....